OpenStudy (gabrielah96):

put the equation into standard form for that particular shape. 4x^2+4y^2-6x+24y=-41

OpenStudy (ranga):

x^2+4y^2-6x+24y=-41 (x^2 - 6x) + (4y^2 + 24y) = -41 complete the square of (x^2 - 6x) and (4y^2 + 24y) first.

OpenStudy (gabrielah96):

its 4x^2

OpenStudy (ranga):

Sorry. 4x^2+4y^2-6x+24y=-41 (4x^2 - 6x) + (4y^2 + 24y) = -41 Complete the square of (4x^2 - 6x) and (4y^2 + 24y) first.

OpenStudy (gabrielah96):

don't i have to take out the 4 first in both?